If it takes Jack 4 days to paint a house and it takes Jill 6 days to paint the same house, how much time would it take Jack and Jill to paint the same house together?
Answer by mangopeeler07(462)

Jack=1/4 of a house a day
Jill=1/6 of a house a day
x=amount of time in days
Together 1/4+1/6 house a day
or 5/12 house a day
5/12x=1
x=12/5 day
or 2 and 2/5 day

Let's translate 2/5 in time
One day=24 hours
2/5 times 24= 9.6

9.6 hours...

.6 hours?

Translate in minutes
.6 of 60 minutes=36 minutes

2 days, 9 hours, and 36 minutes